A man is 25m behind a bus, when the bus stars accelerating at 2ms-2 and the man starts movin with a constant velocity of 10m/s. Time taken by him to board the bus is :
(A) 2s
(B) 3s
(C) 4s
(D) 5s

Grade:11

1 Answers

Mohd huzeifa
5 years ago
Acceleration of man is 0 and acc.  Of bus is 2than relative acc is = 0-2 =-2.     Than 
S = ut -1/2at2
25 =0(t)  -1/2(-2)t2 
t =5 sec.